OpenStudy (vlery):

25^x-14*5^x=-49 find x

OpenStudy (mathmath333):

is it this \(\large\tt \color{black}{25^x-14\times 5^x=-49}\)

OpenStudy (vlery):

yes

OpenStudy (kamille):

think about how 25 can be written in the form of 5?

OpenStudy (kamille):

no ideas?

OpenStudy (vlery):

\[5^{2}\]

OpenStudy (kamille):

correct, so if you have \[25^{x}\] and write in the form of 5,what do you get?

OpenStudy (vlery):

\[5^{2}\]

OpenStudy (kamille):

no, where is x?

OpenStudy (vlery):

\[5^{2x}\]

OpenStudy (mathmath333):

\(\large\tt \color{black}{\text{substitute 5^x=m}}\)

OpenStudy (mathmath333):

\(\large\tt \color{black}{m^2-14\times m=-49}\)

OpenStudy (kamille):

yes,so now you see that you have \[5^{x}\] and \[5^{2x}\] you can notice, that \[5^{x}\] is almost the same as \[5^{2x}\] just one of this is squared,so you can substitute \[5^{x}\] let's say,that \[5^{x}=u\] you get: \[u ^{2}-14u=-49\] can you solve this for me?

OpenStudy (vlery):

couldn't I just put that into the TI calculator and just find the zeros like that?

OpenStudy (anonymous):

u^2 - 14u = -49 u^2 - 14u + 49 = 0 (u -7) (u - 7) = 0

OpenStudy (anonymous):

u = 7 5^x = 7

OpenStudy (vlery):

I don't get it

OpenStudy (kamille):

which part you didn't get?

OpenStudy (kamille):

5^x=7 is not the answer yet

OpenStudy (vlery):

it's just logs that confuse me I'm not sure what formulas to use we just began these thing in class the week

OpenStudy (mathmath333):

use log here

OpenStudy (mathmath333):

take log on both sides

OpenStudy (mathmath333):

\(\large\tt \color{black}{5^x=7}\) \(\large\tt \color{black}{log{5^x}=log7}\) \(\large\tt \color{black}{xlog{5}=log7}\) \(\large\tt \color{black}{x=\dfrac{log7}{log{5}}}\)

OpenStudy (vlery):

I'm getting the tutors help but thank you for helping me